OpenStudy (dan815):

just for fun, see if you think of some ways to prove \[\sum_{k=0}^{m-1}\binom{m-1}{k}=2^{m-1}\]

OpenStudy (empty):

hey bb

OpenStudy (anonymous):

A combinatoric interpretation of this fact is given by counting subsets of different sizes of a set \(S\) of \(m-1\) elements. Since we count the number of subsets of size \(k\) for \(0 \le k \le m-1\), this sum must be equal to the number of subsets of \(S\), which is \(2^{m-1}\).

OpenStudy (anonymous):

Consider \(m-1\) place for a subset. For each subset it can either include or not include an element. For each element, there are \(2\) possibilities. Multiplying these together we get \(2^{m-1}\) subsets.

OpenStudy (anonymous):

You can use binomial theorem.

OpenStudy (anonymous):

As well as power set interpretation

ganeshie8 (ganeshie8):

Just for the sake of alternative, using pascal's rule can be enlightening too \[\binom{n}{k}=\binom{n-1}{k-1}+\binom{n-1}{k} \]  Induction step goes like this \[\begin{align} \sum\binom{m}{k}&=\sum\left[\binom{m-1}{k-1}+\binom{m-1}{k}\right]\\~\\ &=\sum\binom{m-1}{k-1}+\sum\binom{m-1}{k}\\~\\&=2^{m-1}+2^{m-1}\\~\\&=2^{m} \end{align}\]

OpenStudy (amoodarya):

induction over m

OpenStudy (usukidoll):

Binomial Theorem

OpenStudy (amoodarya):

\[A=\left\{ a_1,a_2,a_3,...,a_{m-1} \right\}\\card(a)=m-1\\ \left(\begin{matrix}m-1 \\ 0\end{matrix}\right)+\left(\begin{matrix}m-1 \\ 1\end{matrix}\right)+...+\left(\begin{matrix}m-1 \\ m-1\end{matrix}\right)=total \space \space subsets \space of A\\=2^{card(A)}=2^{m-1} \]